Please answer both parts of the question!!
Answer #1

mass of sample 1 + sample 2 = 100.06 - 50.14 = 49.92 g

density =mass /volume

mass of sample 1 = volume *density = 24.88 * 0.99802 = 24.831 g

mass of sample 2 = 49.92 - 24.831 = 25.089 g

volume of sample 2 = density *mass

= 0.99802 * 25.089 = 25.04 mL

average volume = 25.04 + 24.88 / 2 = 24.96 mL

average deviation = | 24.96 - 24.88 | + | 24.96 - 25.04| / 2

= 0.08+0.08 /2

average deviation = 0.08